The average amount of rain in Mudville is 49 inches per year. The standard deviation is 5 inches. Find the approx. probability that Mudville will receive more than 59 inches in 1 year
Answer by Boreal(15235)

z=(x-mean)/sd=(59-49)/2= 2
probability z > 2 is 0.0228